OpenStudy (anonymous):

Let f(x)=g(x)/h(x) ,g(2)=−1,g ′ (2)=3, h(2)=4,and h ′ (2)=−2.Thenf ′(2)=

OpenStudy (anonymous):

\[f(x) = \frac{g(x)}{h(x)}\]\[f'(x) = \frac{h(x)g'(x) - h'(x)g(x)}{h(x)^{2}}\]\[f'(2) = \frac{(4)(3) - (-2)(-1)}{(4)^{2}} = \frac{10}{16} = \frac{5}{8} = 0.625\]

OpenStudy (anonymous):

Thanks